OpenStudy (anonymous):

Find the integral of: \[\int\limits_{}^{} t \sqrt{4+9t^2 + 16t^4}\]

OpenStudy (anonymous):

\[\int\limits_{}^{} t \sqrt{4+9t^2 + 16t^4}\]

OpenStudy (anonymous):

Not sure if a u-substitution will work, might need to do integration by parts.

OpenStudy (anonymous):

ok, let me try it by parts.

OpenStudy (anonymous):

Alright, so if I was going to do this by parts, how would I split it?

OpenStudy (anonymous):

I would take u=t and dv=the rest : du=dt

OpenStudy (anonymous):

Im thinking \[u = \sqrt{4+9t^2} and dv = \sqrt{16t^4}\]

OpenStudy (anonymous):

doesn't seem right

OpenStudy (anonymous):

No, you can't split up a radical like that.

OpenStudy (anonymous):

How can I integrate that dt then? Doesn't seem to make it easier.

OpenStudy (anonymous):

It looks like along the way during integration-by-parts you can do a u-sub. (or maybe a w-sub, since 'u' has already been used).

OpenStudy (anonymous):

Grr, I just wrote a bunch of equations then my browser crashed.. I'll try going through it another way and let you know if it's any easier.

OpenStudy (anonymous):

ok, thanks